Compute and Analyze Measures for DuPont Disaggregation Analysis
Balance sheets and income statements for Costco Wholesale Corporation follow.

Consolidated Statements of Earnings
For Fiscal Years Ended ($ millions) September 1, 2013 September 2, 2012 August 28, 2011
Revenue
Net Sales $102,870 $97,062 $87,048
Membership fees 2,286 2,075 1,867
Total revenue 105,156 99,137 88,915
Operating expenses
Merchandise costs 91,948 86,823 77,739
Selling, general and administrative 10,104 9,518 8,691
Preopening expenses 51 37 46
Operating Income 3,053 2,759 2,439
Other income (expense)
Interest expense (99) (95) (116)
Interest income and other, net 97 103 60
Income before income taxes 3,051 2,767 2,383
Provision for income taxes 990 1,000 841
Net income including noncontrolling interests 2,061 1,767 1,542
Net income attributable to noncontrolling interests (22) (58) (80)
Net income attributable to Costco $ 2,039 $ 1,709 $ 1,462
Consolidated Balance Sheets
($ millions, except par value and share data) September 1, 2013 September 2, 2012
Assets
Current assets
Cash and cash equivalents $ 4,644 $ 3,528
Short-term investments 1,480 1,326
Receivables, net 1,201 1,026
Merchandise inventories 7,894 7,096
Deferred income taxes and other current assets 621 550
Total current assets 15,840 13,526
Property and equipment
Land 4,409 4,032
Buildings and improvements 11,556 10,879
Equipment and fixtures 4,472 4,261
Construction in progress 585 374
21,022 19,546
Less accumulated depreciation and amortization (7,141) (6,585)
Net property and equipment 13,881 12,961
Other assets 562 653
Total assets $ 30,283 $ 27,140
Liabilities and equity
Current liabilities
Accounts payable $ 7,872 $ 7,303
Accrued salaries and benefits 2,037 1,832
Accrued member rewards 710 661
Accrued sales and other taxes 382 397
Deferred membership fees 1,167 1,101
Other current liabilities 1,089 966
Total current liabilities 13,257 12,260
Long-term debt, excluding current portion 4,998 1,381
Deferred income taxes and other liabilities 1,016 981
Total liabilities 19,271 14,622
Equity
Preferred stock, $0.005 par value:
100,000,000 shares authorized; no shares issued and outstanding 0 0
Common stock, $0.005 par value:
900,000,000 shares authorized;
436,839,000 and 432,350,000 shares issued and outstanding 2 2
Additional paid-in-capital 4,670 4,369
Accumulated other comprehensive (loss) income (122) 156
Retained earnings 6,283 7,834
Total Costco stockholders’ equity 10,833 12,361
Noncontrolling interests 179 157
Total equity 11,012 12,518
Total liabilities and equity $ 30,283 $ 27,140

NOTE: Using net income attributable to controlling interests with your computations, when applicable.

a. Compute ROE for 2013.

Do not round until your final answer. Round answers two decimal places.

ROE =Answer%


b. Confirm that ROE equals ROE computed using the component measures for profit margin, asset turnover, and financial leverage using: ROE = PM * AT * FL.

Compute the components of ROE.

Do not round until your final answer. Round answers two decimal places.

PM = Answer%
AT = Answer
FL = Answer

c. Compute adjusted ROA (assume a tax rate of 37%).

Round answer to two decimal places.
  Adjusted ROA =Answer%
Answer #1

a) ROE = Net Income / Average Shareholders equity

Net Income = $2,039 millions

Average shareholders equity = (Total equity of 2013 + Total Equity of 2012)/2

Average shareholders equity = ($11,012+$12,518)/2 = $11,765 millions.

ROE (2013) = $2,039 millions / $11,765 million × 100 = 17.33%

b) ROE = PM × AT × FL

ROE the ratio into its DuPont

Profit Margin (PM) = Net Income/ Total Revenue × 100

Net Income = $2,039 millions

Total Revenue =$105,516 millions

PM = $2,039/$105,516 ×100 = 1.939% or 1.94%

Assets Turnover (AT) = Total Revenue / Average Total Assets

Total Revenue = $105,516 million

Average total Assets = (Total Assets (2013)+Total Assets (2012))/2

Average total Assets = (($30,283+$27,140)/2)= $28,711.5 millions

Assets Turnover = $105,516millions/ $28,711.5 = 3.6625 or 3.66

Financial Leverage (FL) = Average Total Assets/ Average Total shareholders equity

Average total Assets =$28,711.5 millions

Average total Shareholders equity = $11,765millions

FL = $28,711.5/11,765 millions = 2.4404 or 2.44

ROE= 1.939% × 3.6625× 2.4404= 17.3308% Or 17.33%

c) Computation of Adjusted ROA

Particulars $millions
Income before Income Tax 3,051
Provision of Tax 37% 1,128.87
Net Income Including Non controlling Interest 1,922.13
Net Income attributable to Noncontrolling Interest (22)
Net Income attributable to Costco 1,900.13

Return on Assets = Net Income/ Average Total Assets

Net Income adjusted Tax = $1,900.13 millions

Average total Assets = $28,711.5 millions

ROA = $1,900.13 millions / $28,711.5 millions × 100

ROA = 6.618% Or 6.62%